if Alice deposited $200 in a 2 1/2 year cd at 1.51%, what will the CD make at the end of its term if interest is compounded quarterly. use the compound interest formula and round to the nearest cent.
Answer by lwsshak3(11628)

if Alice deposited $200 in a 2 1/2 year cd at 1.51%, what will the CD make at the end of its term if interest is compounded quarterly. use the compound interest formula and round to the nearest cent.
***
Compound Interest Formula: A=P(1+r/n)^nt, P=initial investment, r=interest rate, n=number of compounding periods per year, A=amt after t-years.
For given problem:
P=$200
r=1.51%
n=4
t=2.5
..
A=200(1+(.015/4))^(4*2.5)
A=200(1+.00375)^(10)
A=200(1.00375)^(10)=207.63
what will the CD make at the end of its term? $207.63